### v1

Abstract
Context It is commonly accepted that the quality of requirements specifications impacts subsequent software engineering activities. However, we still lack empirical evidence to support organizations in deciding whether their requirements are good enough or impede subsequent activities. Objective We aim to contribute empirical evidence to the effect that requirements quality defects have on a software engineering activity that depends on this requirement. Method We replicate a controlled experiment in which 25 participants from industry and university generate domain models from four natural language requirements containing different quality defects. We evaluate the resulting models using both frequentist and Bayesian data analysis. Results Contrary to our expectations, our results show that the use of passive voice only has a minor impact on the resulting domain models. The use of ambiguous pronouns, however, shows a strong effect on various properties of the resulting domain models. Most notably, ambiguous pronouns lead to incorrect associations in domain models. Conclusion Despite being equally advised against by literature and frequentist methods, the Bayesian data analysis shows that the two investigated quality defects have vastly different impacts on software engineering activities and, hence, deserve different levels of attention. Our employed method can be further utilized by researchers to improve reliable, detailed empirical evidence on requirements quality.
Keywords:
Requirements Engineering Requirements Quality Experiment Replication Bayesian Data Analysis
journal: Empirical Software Engineering
1 Introduction
Software requirements specify the needs and constraints stakeholders impose on a desired system. Software requirements specifications (SRS), the explicit manifestation of requirements as an artifact mendez2019artefacts , serve as input for various subsequent software engineering (SE) activities, such as deriving a software architecture, implementing features, or generating test cases mendez2015artefact . As a consequence, the quality of an SRS impacts the quality of requirements-dependent activities frattini2023requirements ; femmer2018requirements ; femmer2015ABREQM . A quality defect in an SRS—for example, an ambiguous formulation—can cause differing interpretations and result in the design and implementation of a solution that does not meet the stakeholders’ needs mendez2017naming . The inherent complexity of natural language (NL), which is most commonly used for specifying requirements franch2017practitioners , aggravates this challenge further. Since quality defects are understood to scale in cost for removal boehm1984software , organizations are interested in identifying and removing these defects as early as possible montgomery2022empirical .
Within the requirements engineering (RE) research domain, the field of requirements quality research aims to meet this challenge montgomery2022empirical . Requirements quality research has already identified several attributes of requirements quality montgomery2022empirical (e.g., unambiguity, completeness, consistency) and proposes quality factors, i.e., requirements writing rules (e.g., the use of passive voice being associated with bad quality femmer2014impact ) as well as tools that automatically detect alleged quality defects femmer2017rapid . However, existing approaches fall short in at least three regards frattini2023requirements : i) only a fraction of publications provide empirical evidence that would demonstrate the impact of quality defects montgomery2022empirical , ii) the few empirical studies that do so largely ignore potentially confounding context factors mund2015does ; juergens2010much , and iii) the analyses conducted in existing publications do not go beyond binary insights (i.e., a quality factor does have an impact or it does notfemmer2014impact ; deissenboeck2007activity . These gaps have impeded the adoption of requirements quality research in practice franch2017practitioners .
In this article, we aim to address the above-mentioned shortcomings by i) replicating the only controlled experiment known to us that contributes empirical evidence on the impact of passive voice on requirements quality femmer2014impact , and strengthening the robustness of this evidence. Further, we ii) collect data about relevant context factors and integrate these data in our data analysis. Finally, we iii) contrast the state-of-the-art frequentist data analysis (FDA) with a Bayesian data analysis (BDA). The latter has recently been popularized in SE research furia2019bayesian since it generates more nuanced empirical insights. Our study is categorized as a laboratory experiment in a contrived setting stol2018abc , isolating the effect of the selected quality factors of interest. The causal inference of their impact contributes to our long-term goal of providing an empirically grounded understanding of the impact of requirements quality. This will support organizations in assessing their requirements and detecting relevant quality defects early.
To that end, we conducted an experiment with 25 participants simulating a requirements-dependent activity (i.e., domain modeling) using four natural-language requirements as input. These requirements contained two common quality defects, passive voice femmer2014impact and ambiguous pronouns ezzini2022automated . Our experiment includes context factors such as experience in software engineering (SE) and RE, domain knowledge, and task experience.
This paper makes the following contributions:
  1. 1.
    a replication of the only controlled experiment investigating the impact of requirements quality;
  2. 2.
    the application of BDA to requirements quality research, which is among the first of its kind in RE.
The remainder of this manuscript is organized as follows. Section 2 introduces relevant related work. We present our research method in Section 3 and the results in Section 4. We discuss these results in Section 5 before concluding our manuscript in Section 6. All supplementary material, including protocols and guidelines for data collection and extraction, the raw data, analysis scripts, figures, and results, are available in our replication package [^1] The replication package is available at https://doi.org/10.5281/zenodo.10423666.
2 Background
Section 2.1 introduces the research domain of this work by summarizing existing research on requirements quality. Section 2.2 motivates BDA—the statistical tool employed in this work—by explaining its adoption in SE research.
2.1 Requirements Quality
Section 2.1.1 introduces the general area of requirements quality research and Section 2.1.2 presents two research directions within. Section 2.1.3 summarizes the three major shortcomings that currently challenge requirements quality research.
2.1.1 Requirements Quality in General
It is commonly accepted that the quality of requirements specifications impacts subsequent SE activities, which depend on these specifications femmer2018requirements ; frattini2023requirements . Quality defects in requirements specifications may, therefore, ultimately cause budget overrun philippo2013requirement or even project failure mendez2017naming . Two further factors aggravate the effect. Firstly, natural language (NL), which is inherently ambiguous and, hence, prone to quality defects, remains the most commonly used syntax to specify requirements franch2023state ; wagner2019status . Secondly, the cost of removing quality defects scales the longer they remain undetected boehm1984software . For example, clarifying an ambiguous requirements specification takes comparatively less effort than re-implementing a faulty implementation based on the ambiguous specification. However, it requires detecting the ambiguity and predicting that the ambiguity potentially causes the implementation to become faulty before it happens. These circumstances necessitate managing the quality of requirements specifications to detect and remove requirements quality defects preemptively.
Requirements quality research seeks answers to this need montgomery2022empirical . One main driver of this research is requirements quality factors frattini2022live , i.e., metrics that can be evaluated on NL requirements specifications to determine quality defects. For example, the voice of a NL sentence (active or passive) is considered a quality factor, as the use of passive voice is associated with bad requirements quality due to potential omission of information femmer2014impact . Automatic detection techniques using natural language processing (NLP) zhao2021natural can automatically evaluate quality factors to detect defects in NL requirements specifications femmer2017rapid .
2.1.2 Existing Research on Passive Voice and Ambiguous Pronouns
We present two examples of commonly researched requirements quality factors in the following sections.
Passive Voice
One commonly researched requirements quality factor is using passive voice in natural language requirements specifications. A sentence in passive voice elevates the semantic patient rather than the semantic agent of the main verb to the grammatical subject ogrady2001linguistics . For example, in the passive voice sentence “Web-based displays of the most current ASPERA-3 data shall be provided for public view.”, the patient of the providing process—the “web-based displays”—becomes the grammatical subject of the sentence. Even though passive voice sentences may still contain the semantic agent (e.g., “Web-based displays of the most current ASPERA-3 data shall be provided for public view by a front-end.”), writers often omit it intentionally or unintentionally krisch2015myth . Figure 1 visualizes the omission of the semantic agent in this exemplary requirement specification.
Figure 1: Formalization of a requirements specification R2 using passive voice
Omitting the semantic agent of a sentence in a passive voice formulation obscures critical information in a requirements specification. Hence, requirements quality guidelines advise against using passive voice pohl2016requirements . However, while several guidelines advise against the use of passive voice based on the theoretical argument of information omission presented above pohl2016requirements ; genova2013framework ; gleich2010ambiguity ; kof2007treatment ; hasso2019detection , only two papers investigate whether passive voice has an actual impact on requirements quality: Krisch et al. let domain experts rate active and passive voice requirements as either problematic or unproblematic. They concluded that most passive voice requirements were unproblematic as the surrounding context information compensated the omission of the semantic agent of the sentence krisch2015myth . Femmer et al. conducted an empirical investigation of the impact of the use of passive voice in requirements specification on the domain modeling activity in a controlled experiment. They concluded that passive voice only causes missing relationships from the domain model, but not missing actors or entities as initially assumed femmer2014impact . The limited evidence for the harmfulness of using passive voice in requirements specifications krisch2015myth ; femmer2014impact stands in stark contrast to the amount of tools and approaches proposed to automatically detect quality defects by identifying the use of passive voice femmer2017rapid ; kof2007treatment ; knauss2009learning ; femmer2018requirements ; hasso2019detection ; rosadini2017using ; soeken2014quality ; drechsler2014automated ; parra2015methodology ; ferrari2018detecting ; genova2013framework .
Ambiguous Pronouns
The inherent ambiguity of natural language poesio1996ambiguity poses several challenges for requirements specifications using natural language nuseibeh2000requirements ; bano2015addressing . One commonly researched requirements quality factor related to ambiguity is the use of ambiguous pronouns, which is a type of referential ambiguity berry2004ambiguity . An ambiguous pronoun exhibits anaphoric ambiguity, that “occurs when a pronoun can plausibly refer to different entities and thus be interpreted differently by different readers” ezzini2022automated . For example, in the requirements specification “The data processing unit stores telemetric data for scientific evaluation; therefore, it needs to comply with the FAIR principles of data storage.”, the pronoun it could syntactically refer to the “data processing unit”, the “telemetric data”, or the “scientific evaluation.” Figure 2 visualizes how a reader can resolve the reference.
Figure 2: Formalization of a requirements specification R3 using an ambiguous pronoun
To avoid deviating interpretations of a requirements specification, established requirements quality guidelines advise against the use of ambiguous pronouns pohl2016requirements at the expense of conciseness. However, the number of publications proposing tools and algorithms to automatically identify and resolve ambiguous pronouns ezzini2022automated ; yang2010extending ; yang2011analysing ; kamsties2000taming ; sharma2016machine ; ezzini2022taphsir ; shah2015resolving ; ezzini2021using ; chantree2006identifying significantly outweighs the singular publication that actually has empirically investigated the effect of ambiguous pronouns. Kamsties et al. investigated the effects of formalizing requirements, which included evaluating the propagation of ambiguous pronouns from NL into more formal specifications kamsties2005empirical . Their experiment involving students revealed that 20-37% of all ambiguous pronouns were incorrectly resolved while formalizing NL requirements specifications. While Kamsties et al. concluded that requirements formalization does not sufficiently resolve ambiguities, these results also support the assumption that ambiguous pronouns propagate into subsequent artifacts depending on the requirements specifications. On the contrary, the scarce empirical work on the effect of ambiguity in general (not specifically ambiguous pronouns) agrees that ambiguity has a negligible effect on downstream software engineering activities de2010ambiguity ; philippo2013requirement . Other than these empirical contributions, the aforementioned publications proposing solutions rather than investigating the relevance of the problem refer to deontic guidelines pohl2016requirements ; belev1989guidelines , anecdotal evidence about ambiguity in general boyd2005measuring ; firesmith2007common ; ferrari2019nlp ; christel1992issues , or—in very rare cases—cognitive science theory poesio1996ambiguity .
2.1.3 Shortcomings in Requirements Quality Research
The previous examples highlight at least three shortcomings from which requirements quality research suffers.
Lack of empirical evidence
First, the relevance of quality factors like passive voice or ambiguous pronouns is rarely determined empirically frattini2023requirements . Scientific contributions proposing solutions (i.e., detecting or removing quality defects) outweigh those investigating the actual extent of the assumed problem. Without knowledge about this extent, it remains unclear whether a proposed solution addresses a problem that is actually relevant to practice.
Previous systematic research has come to the same conclusion. For example, in a previous systematic study, we determined that the effect of quality defects is determined empirically in only 18% of the publications included in our sample frattini2023requirements . Bano et al. found only two publications within their sample of 28 studies that empirically investigated the importance of ambiguity detection bano2015addressing . Montgomery et al. systematically investigated empirical research on requirements quality research and also concluded that most studies focus on improving requirements quality (i.e., detecting and removing defects) rather than defining or evaluating it (i.e., understanding the actual effect) montgomery2022empirical . Instead, most requirements quality publications draw on anecdotal evidence and unproven hypotheses frattini2023requirements . This lack of empirical evidence undermines the trust in requirements quality research and hinders its adoption in practice franch2017practitioners ; phalp2007assessing ; femmer2018scout .
Lack of context
Second, existing research mostly ignores the influence of context factors on the effect of quality defects frattini2023requirements . Context factors encompass all human and organizational factors influencing the downstream SE activities involving requirements specification petersen2009context . For example, the domain experience of a stakeholder or the process model used during development may mediate the effect of ambiguity in requirements specifications philippo2013requirement .
Requirements quality research has acknowledged the relevance of context factors to requirements quality mund2015does ; juergens2010much . Recent propositions have advocated for a shift away from the unrealistic goal of developing a one-size-fits-all solution to requirements quality and, instead, moving towards more context-sensitive research briand2017case ; mendez2017naming . However, this initiative has shown little effect in requirements quality research so far frattini2023requirements .
Lack of detailed projections
Third, the few empirical contributions to requirements quality research limit their insights to categorical projections, i.e., the evaluation of a quality factor on a requirements specification (e.g., using passive voice or not using passive voice) are projected on a categorical scale (e.g., good quality or bad quality). Most commonly, the categorical output space consists of two deissenboeck2007activity (impact or no impact) or three femmer2015ABREQM (positive impact, no impact, or negative impact) categories. This simplification inhibits a nuanced comparison of different quality factors. On an absolute scale, a quality factor having an impact does not automatically entail that this impact is significant and warrants resources for detection and mitigation. On a relative scale, two quality factors that have an impact are impossible to compare to allocate resources towards the more significant one. Consequently, even empirical contributions to the field of requirements quality lack sophisticated insights that would support organizations in determining and dealing with relevant quality factors to control during the RE phase.
2.1.4 Requirements Quality Theory
Based on the identification of the above-mentioned shortcomings, we have developed a requirements quality theory in previous research frattini2023requirements . This theory frames requirements quality as the impact that properties of requirements specifications (called the quality factors) in combination with context factors have on the properties (called attributes) of activities that use these specifications as input. Figure 3 visualizes the requirements quality theory.
Figure 3: Reduced version of the activity-based Requirements Quality Theory frattini2023requirements
The requirements quality theory facilitates overcoming the aforementioned shortcomings. Because the RQT makes the quality of a requirements specification dependent on its impact on subsequent activities, it demands empirical evidence about this impact before claiming that a quality factor reflects actual requirements quality. The inclusion of context factors in the definition of requirements quality mandates context-sensitivity. The abstraction of the impact concept allows for more advanced relationships between specifications and impacted activities than just the categorical type.
However, while the requirements quality research draws on mature software quality research deissenboeck2007activity ; wagner2012quamoco , it has not been actively used yet. Even the predecessor of the theory femmer2015ABREQM was explicitly ignored in follow-up research by its authors due to the complexity of its implementation femmer2017rapid . The work presented in this manuscript constitutes the first application of the theory known to the authors.
2.2 Bayesian data analysis in software engineering
In recent years, SE research has adopted Bayesian data analysis (BDA) for statistical evaluations. BDA signifies a departure from frequentist methods like null-hypothesis significance testing (NHST), the previous state-of-the-art in terms of inferential statistics in SE research. NHST determines whether there is a “statistically significant” difference between two or more distributions. Observations of a dependent variable are stratified by an independent variable to obtain a binary answer of whether or not different values of the independent variable correlate with different distributions of the dependent variable.
Opposed to that, BDA abstains from reducing complex variable distributions to binary inference mcelreath2020statistical . Instead, dependent variables are expressed as a probability distribution, which preserves the natural uncertainty with which any variable is determined. Similarly, the impact of any independent variable on the dependent variable is expressed in terms of a probability distribution. Using Bayes’ Theorem, these assigned prior probability distributions are updated with observed data to obtain a posterior probability distribution furia2019bayesian . Given the observed data, these posterior probability distributions model the most likely impact of variable values. BDA methods are becoming widely adopted also due to the modern computational power enabling Markov Chain Monte Carlo (MCMC) randomized algorithms brooks2011handbook , tools like Stan carpenter2017stan , and libraries like rethinking mcelreath2020statistical and brms burkner2017brms .
While BDA is associated with a much steeper learning curve than frequentist methods, it offers several advantages.
  1. 1.
    BDA is not based on the unsound probabilistic extension of the modus tollens like frequentist hypothesis testing. The modus tollens ($P\rightarrow Q,\neg Q\therefore\neg P$, or if P implies Q and Q is false, then P must also be false) applies to propositional, Boolean logic, but not when inferring from probabilities furia2019bayesian .
  2. 2.
    BDA provides more complex insights than point-wise comparisons. Although BDA lacks out-of-the-box statistical methods like frequentists’ t-tests that are simple to apply, its results reflect the uncertainty of the data, the influence of context, and they can be interpreted more intuitively.
  3. 3.
    BDA makes causal assumptions explicit. The Bayesian workflow gelman2020bayesian makes any hypothesis of causal relations explicit. Analyses become more transparent, and competing causal assumptions are easier to assess.
Furia et al. furia2019bayesian , and Torkar et al. torkar2020missing advocate for the adoption of BDA in software engineering research by discussing its advantages over the frequentist counterpart and mitigating its steep learning curve with extensive demonstrations furia2022applying . SE researchers have begun to apply BDA in various evaluations. Previous studies have used BDA to model bug-fixing time in open source software projects vieira2022bayesian , to confirm the broken window theory in SE leven2022broken , to investigate gender differences in personality traits of software engineers russo2020gender , and to understand data-driven decision making practices svensson2019unfulfilled . In the area of requirements engineering, BDA has been used to evaluate the effect of obsolete requirements on software estimation gren2021possible and to compare requirements prioritization criteria berntsson2021not .
3 Method
We replicate a controlled experiment that investigates the impact of requirements quality on a software engineering activity. Our goal is both to (1) contribute empirical evidence to the effect of quality defects and (2) compare the inferential capabilities of frequentist (FDA) with Bayesian (BDA) statistics. Our experiment is an external, differentiated replication baldassarre2014replication of the study conducted and reported by Femmer et al. femmer2014impact . In this section, we report the design of the replication following the guidelines by Carver carver2010towards .
3.1 Goals
We formulate our goal using the goal-question metric approach wohlin2012experimentation . We aim to characterize the impact of passive sentences and sentences using ambiguous pronouns in requirements on domain modeling with respect to the quality of the created domain model artifacts from the point of view of software engineers in the context of an analysis of requirements from an industrial project. In this definition, software engineer includes all roles that work with requirements specifications, including software developers, requirements engineers, business analysts, managers, and more. We derive the following research questions from our goal:
  • RQ1: Do quality defects in NL requirements specifications harm the domain modeling activity?
    • RQ1.1: Does the use of passive voice in NL requirements specifications harm the duration, completeness, conciseness, and correctness of the domain modeling activity?
    • RQ1.2: Does the use of ambiguous pronouns in NL requirements specifications harm the duration, completeness, conciseness, and correctness of the domain modeling activity?
    • RQ1.3: Does the combined use of passive voice and ambiguous pronouns in NL requirements specifications harm the duration, completeness, conciseness, and correctness of the domain modeling activity?
  • RQ2: Do context factors influence the domain modeling activity?
    • RQ2.1: Do context factors harm the domain modeling activity?
    • RQ2.2: Do context factors mediate the impact of quality defects on the domain modeling activity?
RQ1 is dedicated to the main relationship of interest between quality defects and an affected activity. RQ1.1 directly maps onto the general research question driving the original study femmer2014impact . RQ1.2 extends the scope of the investigation of quality factors with ambiguous pronouns. RQ1.3 investigates the interaction between the two quality factors. RQ2 adds a context-sensitive perspective to the relationship. RQ2.1 focuses on the direct effect that context factors have on the affected activity. RQ2.2 investigates whether context factors mediate the effect of quality defects on the activity.
3.2 Original Experiment
The original study femmer2014impact addresses the research question “Is the use of passive sentences in requirements harmful for domain modeling?” The authors involved 15 university students from different study programs (2 B.Sc., 8 M.Sc., 4 Ph.D., one unknown) in a controlled randomized experiment with a parallel design wohlin2012experimentation . Each participant was assigned to one of two groups and received seven requirements that were formulated either using active or passive voice. The experimental task was to derive a domain model from each requirement that contains all relevant actors, domain objects, and associations between them. The study material and results are available online. [^2] https://doi.org/10.5281/zenodo.7499290
For the dependent variable, the authors calculated the number of missing domain model elements (i.e., actors, objects, and associations). Although the authors also recorded context variables, such as a categorical assessment of general knowledge in SE and RE, these were not used in the analysis. The analysis followed a frequentist approach performing a null-hypothesis significance test for each of the three domain model elements to determine whether a statistically significant difference between the experimental groups exists. The study shows a statistically significant difference in the number of identified associations but not in the number of actors or objects. The authors conclude that the commonly assumed impact of passive voice on missing domain model actors is actually negligible, but passive voice impedes the understanding of the relationships between entities in the requirements specification.
3.3 Our Experiment
We replicate the study by Femmer et al. femmer2014impact because it is the only study that contributes empirical evidence to the impact of passive voice on requirements quality montgomery2022empirical ; frattini2022live in a controlled, experimental setting. During the preparation of our study, we conferred with the authors of the original study. Through this interaction, we identified several opportunities to improve the experimental design. Consequently, we made the following changes to the original study.
  • Experimental design: We employ a factorial crossover instead of a parallel design, which minimizes the risk of confounding (i.e., each participant acts as their own control) while requiring a smaller sample vegas2015crossover .
  • Independent variables: This study investigates—in addition to using passive voice—the impact of ambiguous pronouns in requirements specifications and their combined usage to extend the range of requirements quality defects.
  • Dependent variables: We merged two types of elements in the domain model (the nodes of the model, i.e., actors and objects) into a single type entity. This decision was taken after feedback from a pilot study. Furthermore, we increased the dependent variables by additionally evaluating the number of superfluous entities, the number of wrong associations, and the duration for creating the domain model.
  • Sampling strategy: We sample from both students and practitioners of software engineering to more accurately represent the target population of software engineers. This change aims at increasing the external validity of our results baltes2022sampling .
  • Instrumentation: The participants performed the experimental task online using a web-based application rather than offline using pen and paper. This allows for more flexibility in reaching industry participants baltes2022sampling .
  • Context factors: To obtain a richer understanding of the impact of quality defects, we included seven additional context factors. This change made it necessary to extend the questionnaire used in the original study to collect demographic information from the participants.
  • Experimental object: We sampled the objects from a data set of industrial requirements specifications ferrari2017pure rather than from a requirements specification written in a student project femmer2014impact to increase the realism of the experimental task sjoberg2003challenges .
  • Analysis: The crossover design produces paired data as opposed to the unpaired data of the original experiment, which changes the appropriate hypothesis test vegas2015crossover (t-test in the original study vs. Wilcoxon rank sum test in this study). In addition, we extend the original FDA by performing a Bonferroni correction to deal with family-wise error rate when testing multiple hypotheses benjamini1995controlling . Furthermore, we additionally analyze the data using BDA.
In the rest of this subsection, we report the design of our experiment following the guidelines by Jedlitschka et al. jedlitschka2008reporting .
3.3.1 Experimental Task
We simulate the use of a requirements specification by subjecting participants to a requirements processing activity, i.e., a common task representing the use of requirements femmer2014impact . In particular, we present four single-sentence, natural language requirements to the participants and request them to derive a domain model for each of them. Figure 4 visualizes the expected domain model for the requirement “Every research object is represented in a JSON-LD format and stored in a document database if it contains a CC license.” which contains both of the two seeded quality defects. These defects result in the following challenges according to literature pohl2016requirements :
  1. 1.
    The verb in passive voice omits an important entity of the requirement; i.e., that the data processing unit stores the research object in a document database (Label 1 in Figure 4).
  2. 2.
    The ambiguous pronoun “it” can syntactically be connected to several preceding noun phrases (“Every research object”, “JSON-LD format”, and “a document database” or the implicit “Data Processing Unit”) by a reader but semantically only applies to the research object (Label 2 in Figure 4).
Figure 4: Domain modeling task example for requirement 4.
The goal of the experimental task is to derive a semantically correct domain model from the natural language requirement which includes identifying all entities (including the implicit ones) and connecting these entities correctly (including those derived from syntactically vague associations).
The selection of dependent variables was driven by the activity-based requirements quality theory femmer2015ABREQM ; frattini2023requirements . Accordingly, requirements quality is measured by the effect that quality factors have on the relevant attributes of requirements-dependent activity. We selected the following dependent variables representing the relevant attributes of the domain-modeling activity with the given motivation:
  • Duration: the longer the domain modeling task takes, the more expensive it is.
  • Number of missing entities: entities missing from the domain model produce potential cost for failing to involve the respective actor or object.
  • Number of superfluous entities: entities added to the domain model but not implied by the requirement unnecessarily constrain the solution space.
  • Number of missing associations: associations missing from the domain model produce a potential cost for failing to identify a dependency between two entities.
  • Number of wrong associations: associations connecting two entities that establish an unnecessary dependency between them while neglecting an actual dependency.
We characterize the domain modeling activity in terms of immediacy (duration), completeness (missing entities and associations), conciseness (superfluous entities), and correctness (wrong associations).
3.3.2 Hypotheses
[…]
“There is no difference in {dep} of the domain models based on requirements specifications containing no quality defect and requirements specifications containing {ind}.”
To capture the context of the experiment, we collected factors based on related work petersen2009context ; mund2015does ; juergens2010much , including the experience of a practitioner regarding software and requirements engineering, but also in SE roles and in the modeling task itself. Additionally, we assume that the practitioners’ education and domain knowledge influence the dependent variables. Table 1 summarizes the variables involved in this study.
Table 1: Variables of the study (independent, context, and dependent).
Variable Name Type Description Data type Range
Requirements Quality Defect
RQD
ind
The use of a verb in passive voice, an ambiguous pronoun, or both
categorical
{none, PV, AP, PVAP}
Experience in SE
exp.se
con
Years of experience in software engineering
count
$\mathbb{N}$
Experience in RE
exp.re
con
Years of experience in requirements engineering
count
$\mathbb{N}$
Education
edu
con
Highest acquired degree
ordinal
{High School, B.Sc., M.Sc., Ph.D.}
Primary role
role
con
SE-related Role with the most years of professional experience
categorical
{requirements engineer, product owner, software architect, developer, tester, quality engineer, trainer, manager, none}
Task experience
exp.task
con
Experience with the task of domain modeling
ordinal
{never, rarely, from time to time, often}
Formal modeling training
formal
con
Formal training in in domain modeling
categorical
{true,false}
Domain knowledge in {domain}
dom. {domain}
con
[…]
ordinal
{1,2,3,4,5}
Duration
$D$
dep
Number of minutes it took the participant to complete the experimental task on one requirement
count
$\mathbb{N}$
Missing entities
$E^{-}$
dep
Number of relevant entities missing from the submitted domain model
count
$[0,E_{expected}]$
Superfluous entities
$E^{+}$
dep
Number of not relevant entities included in the submitted domain model
count
$\mathbb{N}$
Missing associations
$A^{-}$
dep
Number of associations missing from the submitted domain model
count
$[0,A_{expected}]$
Wrong associations
$A^{\times}$
dep
Number of associations where either the source or target of the edge is a different entity than implied by the requirement
count
$[0,A_{found}]$
3.3.3 Experimental Design
Our experimental design includes one factor (RQD) representing the alleged quality defect seeded in a requirements specification. This main factor contains four treatments: a control one (no defects) and three experimental ones (passive voice (PV), ambiguous pronoun (AP), and both (PVAP)).
Given our sampling strategy involving industry practitioners, which are difficult to recruit for controlled experiments sjoberg2003challenges , we anticipated a moderate sample size of participants. Consequently, we opted for a crossover design instead of a parallel design, i.e., we apply every treatment to all subjects instead of distributing the subjects among the treatments. Previously, Kitchenham et al. advised against the use of crossover designs kitchenham2003case . Mainly, the validity of crossover design experiments is challenged by the following confounding factors:
  1. 1.
    the period in which a treatment is applied to a subject, as certain periods may influence the dependent variables (e.g., measuring the dependent variable in the period right before an exam, participating students may be more tense and less focused);
  2. 2.
    the sequence, in which the treatments are applied to a subject, as certain sequences may entail a learning effect that systematically influences a dependent variable (e.g., repeatedly performing a task might improve the performance of experiment participants); and
  3. 3.
    the previous treatment (also called the carryover effect, as the application of a treatment may have an effect on the dependent variable when applying a subsequent treatment vegas2015crossover .
However, recent adoptions of best practices from other disciplines made this design applicable to SE research without compromising the validity of the results vegas2015crossover ; fucci2016an . The threats to validity can be mitigated during design and analysis vegas2015crossover by (1) randomizing the order of treatments and (2) including the independent variables’ period, sequence, and the interaction between them in the analysis. When controlling the threats to validity, the crossover design provides two benefits. Firstly, it requires fewer participants, as an experiment with $n_{p}$ participants and $n_{t}$ treatments yields $n_{p}\times n_{t}$ observations instead of only $n_{p}$ wohlin2012experimentation . Secondly, it accounts for differences between participants, as the dependent variables can be measured in relation to the average response of each subject instead of the average response of each treatment group kitchenham2003case . Therefore, each subject acts as its own control and mitigates within-subject variance.
Each experimental session contained four main periods in which we applied one treatment to the subject. We randomized the order of treatment application to disperse the confounding sequence and carryover effect brown1980crossover ; vegas2015crossover . This resulted in 24 unique sequences of treatment application ($n_{t}!=24$) and, consequently, 24 experimental groups. The experiment was single-blinded—i.e., the participants did not know the requirements’ sequence, but the researchers did.
3.3.4 Objects
The experimental object consisted of four English, single-sentence NL requirements specifications $R_{1}$-$R_{4}$. An additional warm-up object ($R_{0}$) preceded the actual experimental objects, adding a fifth experimental period to each session. It was only used to familiarize the participants with the experimental task and tool and was not considered in the data analysis. The four experimental objects were manually seeded with defects corresponding to our four treatments: one requirement containing none of the two faults, one containing a verb in passive voice, one containing an ambiguous pronoun, and one containing both a verb in passive voice and an ambiguous pronoun. The requirements’ mean length is 17.8 words (sd=4).
The first author derived the experimental objects from the requirements specification of the Mars Express ASPERA-3 Processing and Archiving Facility (APAF), a real-world specification from the PuRE data set ferrari2017pure . From this requirements specification, the first author selected five single-sentence natural language requirements and modified them to ensure two defect-free requirements (one warm-up object $R_{0}$ and one for the defect-free baseline $R_{1}$) and three objects with the respective defects ($R_{2}$-$R_{4}$). The second author reviewed and adjusted the selected objects.
3.3.5 Subjects
The target population of interest consists of people involved in software engineering which work with requirements specifications. We used a non-probability sampling approach based on a mix of purposive and convenience sampling baltes2022sampling . In particular, we wanted to select participants who are diverse in terms of the context variables as determined in Table 1, including their experience, education, and software engineering roles. We approached both students participating in RE courses at our respective institutions and practitioners in our collaborators’ network to purposefully diversify the experience and education of our sample. For all other demographic context factors (e.g., SE roles) we had to rely on convenience sampling.
We approached 52 potential candidates (32 practitioners and 20 students), and 27 candidates (19 & 8) agreed to participate in the experiment (response rate of 52%). Two students did not show up to the agreed time slot. Our final sample includes 19 practitioners from six different companies and six students from two universities. Participation in the experiment was entirely voluntary and not compensated. The number of participants ($n_{p}=25$) exceeded the number of experimental sequences ($n_{t}!=24$) such that we had at least one subject per experimental group and evenly dispersed any confounding sequence or carryover effect vegas2015crossover .
Figure 5 shows the distribution of experiment subjects’ experience in SE and RE in years. which is widespread in our sample. Among the 25 participants, the reported primary roles are developer (10), architect (6), requirements engineer (5), manager (1), and no prior professional role (3). Students who have not yet had a professional software engineering role constitute this last group. The distribution covers many SE-relevant roles but excludes others, such as testers or product owners.
Figure 5: Distribution of SE and RE experience.
Among the $25$ participants, $5$ reported a high school degree as their level of education, $8$ a Bachelor’s degree, and $12$ a Master’s degree. No participant reported a Ph.D. degree as their highest degree of education. Figure 6 visualizes the distribution of the participant’s experience in the four domains that were relevant to the experimental requirements: aeronautics, telemetry, databases, and open source. For the latter two, results are balanced across the different knowledge levels, while the former two confirm our assumption that all participants had a low-level knowledge of aeronautics and telemetry systems.
Figure 6: Distribution of knowledge in the four relevant domains.
The experiment tool—the Google Draw plugin within the Google Document—was unknown to most ($18$ never used it, $6$ rarely, and $1$ from time to time). A total of $16$ participants ($64$%) report having received a form of training in the modeling activity. The modeling experience (never: $1$, rarely: $11$, from time to time: $10$, often: $3$) resembles a normal distribution.
Given the distribution of responses in these context variables, we disqualified the following predictors: aeronautics domain knowledge, telemetry domain knowledge, and experience with the experiment tool. These variables are not sufficiently distributed in our sample of study participants, i.e., several categories of these variables are underrepresented. Consequently, they are unable to effectively block the influence of that variable on the dependent variable wohlin2012experimentation .
3.3.6 Instrumentation
We used a Google Docs document [^3] https://www.google.de/intl/en/docs/about/ for the task and a Google Form questionnaire [^4] https://www.google.com/forms/about/ to collect demographic information. The Google Docs document lent itself to the task due to its accessibility and its simple modeling tool with the embedded Google Drawings. The modeling tool represented the optimal trade-off between complexity—as neither previous knowledge nor additional software was necessary to conduct the experimental task—and suitability—as it contains all elements relevant to the domain modeling task (i.e., nodes for entities and edges for associations). This main study document explained the experimental task, an example of the domain modeling task, and a short context description of the system from its original requirements specification ferrari2017pure .
We created a survey questionnaire to collect demographic information relevant to the experiment using Google Forms. All participants could answer the survey only after completing the task to avoid fatiguing effects. At the beginning of the questionnaire, participants entered their assigned participant ID (PID), such that we could connect their response to the experimental task to their response to the questionnaire without storing any personal data. The questions were designed to collect all relevant independent variables listed in Table 1.
We piloted the experiment in a session with two Ph.D. students in SE. We clarified the instruction text and task descriptions based on the collected feedback.
3.3.7 Data Collection Procedure
We scheduled a one-hour session according to the availability of the participants. Because of differing schedules and time zones, we scheduled 16 sessions with up to three participants simultaneously. We conducted the sessions between 2023-04-03 and 2023-04-17.
Each session started with the first author explaining the general procedure of the experiment and obtaining consent to evaluate and disclose the anonymized data. No participant refused this consent and all data points could be included in the data evaluation procedure. Then, participants were instructed to read the prepared document in order and complete the contained tasks. The document contained all descriptions about the task such that all participants received the same instructions. The first author oversaw all sessions to address technical difficulties, and recorded the minutes each participant spent per period. Ten minutes were estimated per period, but participants were free to allocate their time. In case participants took longer than the scheduled one hour, they completed the task in as much time as they required. Once the task was complete, participants also filled in the questionnaire to provide demographic information on context variables.
3.3.8 Data Preparation
To evaluate the collected data, we created a code book that characterizes issues in domain models. We developed detection rules for each dependent variable of the resulting product (i.e., missing entity, superfluous entity, missing association, and wrong association). Then, the first author manually evaluated the resulting domain models of each participant using this codebook and recorded all detected issues.
The result of the coding process was a table where one row represents the evaluation of one domain model. Given $n_{p}=25$ participants and $n_{r}=4$ requirements, we ended up with $n_{p}\times n_{r}=100$ data points. Each data point contained the number of issues of each of the four types that occurred in the respective domain model. Finally, we standardized numerical variables in the demographic data for easier processing.
To assess the reliability of the rating, the fourth author of the paper independently recorded issues of three randomly selected participant responses, yielding an overlap of twelve ratings. Since each domain model can contain an arbitrary number of issues of each type, but our dependent variables only model the number of times that an issue type occurred, we consider each rating of a domain model as a vector of dimensions equal to the number of issue types. We then calculated the agreement between two ratings of the same domain model using the cosine similarity between the vectors. The average cosine similarity is 87.23% and represents substantial agreement.
3.3.9 Frequentist Data Analysis
We performed a frequentist data analysis of the experimental data as in the original experiment femmer2014impact . Since the factor is categorical and all dependent variables are continuous, our statistical method of choice falls between the parametric paired t-test hsu2014paired or the non-parametric Wilcoxon rank sum test wilcoxon1945individual based on the distribution of the variables, which we evaluated using the Shapiro-Wilk test results shapiro1965analysis .
We reject a null hypothesis if the resulting p-value of a statistical test is lower than the significance level $\alpha$. To account for type I errors when performing multiple hypotheses tests targeting the same independent variable, we apply the Bonferroni correction benjamini1995controlling : we considered $\alpha^{\prime}=\frac{\alpha}{m}$ where $\alpha=0.05$ and $m$ is the number of hypotheses tested for each value of the independent variable. For our five families of hypotheses $\alpha^{\prime}=\frac{0.05}{5}=0.01$.
Additionally, we report the effect size dybaa2006systematic using Cohen’s D for the paired student t-test cohen1969statistical and the matched-pairs rank biserial correlation coefficient for Wilcoxon rank sum test king2018statistical .
3.3.10 Bayesian Data Analysis
We apply Bayesian data analysis with Pearl’s framework for causal inference pearl2016causal to complement the frequentist data analysis furia2019bayesian . Given the limited adoption of Bayesian data analysis in software engineering research furia2019bayesian ; furia2022applying ; furia2023towards , we complement this method section with a running example for understandability. In this running example, we illustrate the methodological steps of Bayesian data analysis for the hypothesis that requirements quality defects influence the number of wrong associations in a resulting domain model.
Three steps siebert2023applications , which are the major steps of Pearl’s original model of causal statistical inference pearl2016causal , comprise the analysis. We explain each step in the following paragraphs.
Modeling
In the modeling step, we make our causal assumptions about the underlying effect explicit in a graphical causal model. The graphical causal model takes the form of a directed acyclic graph (DAG) in which nodes represent variables and directed edges represent causal effects elwert2013graphical . Our DAG contains four groups of variables:
  1. 1.
    Treatment: The independent variable, which represents the requirements quality defect present in the requirement.
  2. 2.
    Context factors: The independent variables, which represent the properties of the participants.
  3. 3.
    Confounding factors: The independent variables, which represent the crossover experiment design (e.g., session, period).
  4. 4.
    Response variables: The dependent variables.
The effect of the treatments on the response variables is the subject of the analysis. By including both context and confounding factors, their influence is factored out from the treatments’ causal effect on the response variables. Consequently, the effect of interest can be isolated from any confounding factor included in the DAG. We assume causal relations—represented by edges in the DAG—between every independent (treatment, context, and confounding) and the dependent variables. Additionally, independent variables may influence other independent variables.
Figure 7 shows the DAG of the running example. The treatment, response variable, and context factors correspond to the study variables as outlined in Table 1. The confounding factors represent the influence of the experimental design. The experimental period blocks the learning effect, i.e., the influence on the response variable caused by repeatedly performing the task. The duration blocks the time effect, i.e., the influence on the response variable caused by the amount of time that a participant took for each instance of the task.
Figure 7: DAG for the analysis of wrong associations
The DAG does not visualize the interaction effects we assume between two independent variables. An interaction effect occurs when the influence of one independent variable on the dependent variable depends on the value of another independent variable mcelreath2020statistical . Visualizations of interaction effects in DAGs have been proposed nilsson2021directed but are not common practice. In the running example, we assume two interaction effects via the following hypotheses:
  1. 1.
    requirements quality * domain knowledge: domain knowledge can compensate the effect of requirements quality defects poesio1996ambiguity
  2. 2.
    requirements quality * period carryover effect vegas2015crossover : the effect of a treatment may be influenced by the treatments applied in previous periods
Identification
Including an independent variable that has an association with another independent variable may introduce a previously blocked confounding effect by opening a backdoor pearl1995bayesian (introducing spurious associations). Consequently, blindly moving forward with all variables may harm the causal analysis. Instead, the so-called adjustment set of variables needs to be selected mcelreath2020statistical in the identification step. A series of four criteria mcelreath2020statistical allows to make an informed selection of variables to include in the final estimation step. This way, we avoid variable bias like colliders which confound the causal effect between the treatment and the response variable.
In the running example, we assume the following causal relation between independent variables. The more experience a participant has in SE or RE, the more likely it is that they have acquired respective domain knowledge (experience in SE/RE $\rightarrow$ domain knowledge). We need to consider this relationship in the next step to avoid attributing impact to the wrong independent variable. For instance, in the running example, we need to distinguish whether experience in SE/RE has a direct influence on wrong association or whether it just has an influence on domain knowledge, which has an influence on the response variable.
Because we employ an experiment as our research method and fully control the treatment, there is no influence of any other independent variable on the treatment variable.
Estimation
In the estimation step, we perform a regression analysis. The regression analysis results in estimates of the response variable depending on the values of the independent variables. The result of the regression analysis is a Bayesian model trained with empirical data. The model provides the magnitude and sign of the effect that each independent variable has on the dependent response variable.
The estimation step begins by selecting a distribution type (likelihood) that represents the dependent response variable gelman2020bayesian . We select the distribution type based on the maximum entropy criterion jaynes03 and ontological assumptions. This means we select the least restrictive distribution that fulfills all ontological assumptions about the variables’ properties.
In our running example, the response variable is a count of wrong associations in a domain model. Consequently, the distribution must be discrete and only allow positive numbers or zero. Additionally, the response variable is bounded by the number of expected associations of the domain model, i.e., the number of associations in the sample solution, since a participant can only connect as many associations wrongly in the model as there were associations expected. Any associations added beyond the expected associations count as superfluous associations, a different response variable. Consequently, we represent the response variable with a Binomial distribution. The following formula encodes that the number of wrong associations in one domain model i ($E_{i}^{\times}$) is distributed as a Binomial distribution with the number of trials equal to the number of expected associations ($E$) and a probability $p_{i}\in[0,1]$ of getting one association wrong.
$E_{i}^{\times}\sim Binomial(E,p_{i})$
This formula assumes that the event—connecting one association wrong—is independent, i.e., one wrong association does not influence the success of any other association.
In the next step, we define the parameter that determines the response variable distribution (in the running example: $p_{i}$) in relation to the predictors selected in the identification step. The following formula shows a simplified version of this parameter definition (excluding most of the previously mentioned predictors in Table 1 for brevity).
$logit(p_{i})=\alpha+\alpha_{PID}+\beta_{RQD}^{T}\times RQD_{i}+\beta_{SE}\times exp.se_{i}$
The logit operator scales the parameter $p_{i}$ to a range of $[0,1]$ since the probability parameter of the Binomial distribution only accepts this range of values demaris1992logit ; mcelreath2020statistical . The parameter $p_{i}$ is, in this example, determined by the following predictors:
  1. 1.
    Intercept ($\alpha$): the grand mean of connecting an association wrongly, i.e., the baseline challenge of getting an association wrong.
  2. 2.
    Group-level intercept ($\alpha_{PID}$, where the results of one participant represent one group): the participant-specific mean of connecting an association wrongly, i.e., the within-subject variance of response variables vegas2015crossover modeled via partial pooling ernst2018bayesian
  3. 3.
    Treatment ($RQD_{i}$): the influence of a requirements quality defect on the probability of connecting an association wrong (as an offset from the grand mean).
  4. 4.
    Software Engineering Experience ($exp.se_{i}$): the influence of the subject’s software engineering experience on the probability of connecting an association wrong (as an offset from the grand mean).
The variables $RQD_{i}$ and $exp.se_{i}$ contain the values recorded during instance $i$ of conducting the experimental task and are each prefixed with coefficients $\beta_{RQD}^{T}$ and $\beta_{SE}$. These coefficients are Gaussian probability distributions that represent the magnitude and direction of the influence that the variable values have on the parameter $p_{i}$ and, therefore, on the distribution of the response variable. The mean $\mu$ of the coefficient represents the average effect of the variable on the parameter $p_{i}$, and the standard deviation $\sigma$ encodes the variation around this average effect. A standard deviation of $\sigma=0$ would mean that the variable has a deterministic effect of strength $\mu$ on $p_{i}$ and, therefore, the distribution of the response variable. In reality, this is highly unrealistic. Hence, the standard deviation captures the uncertainty of the effect of a variable on $p_{i}$.
Special cases of variables are the intercepts $\alpha$ and $\alpha_{PID}$, which are probability distributions without any variable and, hence, represent the predictor-independent general and participant-specific probability of connecting an association wrong. In the beginning, we assign probability distributions spread around $\mu=0$ ($\beta_{RQD}^{T}\sim Normal(0,0.5)$), so-called uninformative priors, to these coefficients. These distributions encode our prior beliefs about the influence of the respective predictor, i.e., that it is yet unknown whether the predictor has a positive ($\mu>0$) or negative ($\mu<0$) influence on $p_{i}$. Only where previous evidence for the impact of a predictor on the response variable exists, we select more informative priors. For example, the experiment by Femmer et al. femmer2014impact indicates that missing entities and missing associations are in general rare, which we represent in our priors by selecting $\alpha\sim Normal(-1,0.5)$ for the intercept.
We assess the feasibility of the selected prior distributions via prior predictive checks wesner2021choosing . During this check, we sample only from the priors, i.e., we predict the response variable given the recorded data of independent variables and the prior probability distributions of the predictor coefficients. Figure 8 visualizes the result of the prior predictive check. The grey bars represent the actual observed distribution of the response variable. For example, 75 domain models contained zero wrong associations. The distribution of predicted values for the response variable (cyan whisker plots) encompasses the actual observed distribution of the response variable. This confirms that the actually observed distribution is approximately determined by the uninformative prior distributions.
Figure 8: Predictive checks with prior and updated posterior coefficient distributions
Upon confirmation of the priors’ feasibility, we train the Bayesian models with the data recorded during the experiment. We conducted the analysis using the brms library burkner2017brms in R. Hamiltonian Monte Carlo Markov Chains (MCMC) brooks2011handbook update the coefficient distributions based on the empirical data. During this process, the parameters of the coefficient distributions are adjusted to better reflect the response variable based on the predictor variable values.
After the training process, we perform posterior predictive checks, which work similarly to the prior predictive check but use the updated posterior coefficient distributions instead of the prior distributions. Figure 8 also visualizes the posterior predictive check for the running example. The distribution of the predicted values (red whisker plots) still encompasses the actually observed distribution of the response variable but has narrowed around these values. This indicates that the posterior distributions encode the influence of the predictor variables more accurately than the prior distributions, i.e., that the model has successfully gained predictive power during the training process.
To overcome the problem mentioned in the identification step, i.e., attributing impact to the wrong predictor, we train additional models per response variable to test for conditional independence mcelreath2020statistical . For example, to determine the correct causal relationship between the two independent variables experience in SE, domain knowledge, and the dependent variable, we train two additional models where each one misses one of the two variables furia2023towards . After training, we compare the posterior distributions of the remaining parameter coefficients. If a posterior distribution significantly moves from $|\mu|>0$ towards $\mu=0$ when including a variable, then the response variable is independent of that variable when conditioning on the included variable. The model does not gain any further information from the variable with $\mu\simeq 0$, and its causal relation is disputed. If the posterior distribution does not deviate significantly when including another variable, its causal impact is confirmed.
Finally, we perform a stratified posterior prediction to answer our research questions. To this end, we construct a synthetic data set with four data points, one for each value of the main factor variable (i.e., baseline, PV, AP, PVAP). We fix all other independent variables at representative values—i.e., the mean for continuous and the mode for discrete variables. Then, we sampled $6,000$ predictions for each of the four data points. This isolates the effect of the treatment but maintains the uncertainty of the influence of every independent variable encoded in the standard deviation of every predictor coefficient and, hence, more accurately describes the causal relationship between the treatment and the outcome. We compare the $6,000$ predictions of each of the three treatments (PV, AP, PVAP) with the $6,000$ predictions from the baseline (no defect) and count how often the treatment causes a higher, equal, or lower outcome variable. We scale these values to percentages to summarize the effect of the treatment on the outcome variable. This evaluation avoids a point-wise reduction of the results and comparison to an arbitrary significance level as customary in frequentist analyses gren2021possible . Consequently, we abstain from a binary answer to the hypotheses and rather present the more informative distribution of results.
Additionally, we plot the marginal effect of selected independent variables to visualize their isolated impact on the response variable. The isolated impact reveals how context and confounding factors influence the response variable. This includes visualizing the carryover effect, i.e., the interaction between the treatment and the period.
4 Results
Section 4.1 shows the results of the frequentist and Section 4.2 the results of the Bayesian data analysis. Section 4.3 compares our results to the original study.
4.1 Frequentist Data Analysis
Table 2 lists the results of our frequentist data analysis and relates them to the results from the original study femmer2014impact .
Outcome Treatment Original femmer2014impact Replication
$p$ CI ES $p$ CI ES
Duration PV 0.67 $(-0.4,0.7)$0.13
AP 0.86 $(-0.7,0.86)$ 0.01
PVAP 0.49 $(-0.5,0.25)$ 0.14
Missing Actors
PV 0.10 $(0,\infty)$ 0.39
Missing Entities PV 0.25 $(-1,\infty)$ 0.25 $\ll$ 0.01 $(-1,0)$0.79
AP $\ll$ 0.01 $(-1.5,0)$0.93
PVAP $\ll$ 0.01 $(-2,0)$0.81
Superfluous Entities PV 0.64 $(-0.5,0)$ 0.14
AP 0.19 $(-2.5,0)$0.41
PVAP 0.62 $(-1,1)$ 0.15
Missing Associations PV 0.02 $(1,\infty)$ 0.75 0.025 $(-1,0)$0.58
AP $\ll$ 0.01 $(-2,-1.5)$0.87
PVAP $\ll$ 0.01 $(-2,-1)$0.84
Wrong Associations PV 1.0 $(0,0)$ 0.0
AP $\ll$ 0.01 $(-1,0)$0.85
PVAP 0.052 $(-1.5,0)$0.67
Table 2: Results of frequentist analysis including the p-value of the hypothesis test (p), confidence interval (CI), and effect size (ES). Statistically significant results in bold (original study $\alpha=0.05$, this replication $\alpha^{\prime}=0.01$).
The frequentist data analysis suggests rejecting the following hypotheses and, therefore, proposes the following effects as statistically significant (with $\alpha^{\prime}=0.01$):
  1. 1.
    $H_{0}^{PV\rightarrow E^{-}}$: passive voice impacts the number of missing entities
  2. 2.
    $H_{0}^{AP\rightarrow E^{-}}$: ambiguous pronouns impacts the number of missing entities
  3. 3.
[…]
  4. 4.
    $H_{0}^{AP\rightarrow A^{-}}$: ambiguous pronouns impact the number of missing associations
  5. 5.
[…]
  6. 6.
    $H_{0}^{AP\rightarrow A^{\times}}$: ambiguous pronouns impact the number of wrong associations
The associated effect size is considered large cohen2008explaining in all cases.
4.2 Bayesian Data Analysis
This section follows the methodology described in Section 3.3.10 by presenting the DAG in Section 4.2.1, posterior predictions in Section 4.2.2, and marginal plots in Section 4.2.3.
4.2.1 Causal Model and Adjustment Set
Figure 9 visualizes the DAG, that graphically models our causal assumptions. It is an extension of Figure 7, the running example, including all five dependent variables. To preserve readability of the DAG, we introduce a distributor node. This node substitutes the connections from the source of every incoming edge to the target of every outgoing edge.
Figure 9: Directed acyclic graph visualizing all causal assumptions
The edges represent the same causal reasoning as presented in the running example in Section 3.3.10. We assume that all independent variables (treatments, context factors, and confounding factors) have an impact on all dependent variables. The impact of the requirements quality defect (the three treatments) is the relationship of interest in our analyses. Additionally, to the already described impact of experience in SE on domain knowledge, we assume the following causal relationships:
  1. 1.
    Duration impacts all other dependent variables: Since we did not constrain the time for each period, different amounts of minutes taken for each object may influence the results. Taking a longer time for one domain model may reduce the amount of defects in the final model.
  2. 2.
    Missing entities impact missing associations: If an entity is missing from the domain model, any association involving that entity will also be missing.
In the second step in the statistical causal inference siebert2023applications , the identification step, we found the adjustment set to include all variables for prediction. To discern the impact of independent variables with causal relations among them, we developed comparison models in which certain variables were excluded. The comparison showed that the exclusion did not change the estimations of the coefficient distributions. Hence, we assume all causal relationships are feasible and evaluate the full model, including all eligible variables as predictors.
4.2.2 Posterior Predictions
Based on maximum entropy jaynes03 , we model the five response variables using the following probability distributions. The duration is centered around the global mean, therefore, we model it with a Gaussian distribution around $\mu=0$. The number of superfluous entities is an unbounded count with an index of dispersion of about 1.5, hence, we model it as a negative binomial distribution. Missing entities, missing associations, and wrong associations are bounded counts and, hence, modeled as Binomial distributions. Table 3 contains the result of the predictions from the posterior distributions. Each cell contains the resulting likelihood that the occurrence of a factor causes fewer or more issues of the respective outcome compared to the baseline of no quality defects [^5] The remaining cases ($100\%-less-more$) are omitted from the table. The larger the difference between the likelihood of more ($+$) than fewer (--) issues, the stronger the effect of that factor on the outcome. If the likelihood of more ($+$) issues outweighs the likelihood of fewer (--) issues, the factor has a negative effect. If the two values are similar, then the factor has no clear effect on the outcome.
Outcome PV AP PVAP
Duration 57.2% 42.8% 51.7% 48.3% 44.8% 55.2%
Missing entities 29.6% 32.5% 24.7% 40.1% 27.4% 35.6%
Superfluous entities 26.6% 22.5% 22.5% 33.3% 26.4% 24.6%
Missing associations 25.0% 45.0% 20.2% 51.2% 22.2% 49.4%
Wrong associations 10.5% 11.5% 5.2% 44.6% 6.9% 31.5%
Table 3: Likelihood that a treatment produces fewer (--) or more ($+$) occurrences of the respective outcome variable.
For example, the first two cells in the second row of Table 3 state that using passive voice causes fewer missing entities in $29.6\%$ and more missing entities in $31.5\%$ of all cases. In the remaining $38.9\%$ of all cases, passive voice causes neither more nor fewer missing entities. Given this balance, the effect of passive voice on missing entities is unclear, and there is not enough evidence to reject $H_{0}^{PV\rightarrow E^{-}}$.
[…]
4.2.3 Marginal and Conditional Effects
Marginal plots visualize the isolated effect of specific predictors when fixing all other predictors to representative values. In the following, we present selected marginal plots that show the effects of interest. The remaining plots can be found in our replication package.
Missing entities impact missing associations
Figure 10 visualizes the effect of the number of missing entities on the number of missing associations. The $y$-axis represents the expected value of missing entities over multiple attempts with a trial size of one. Hence, it corresponds to the likelihood of missing one entity.
Figure 10: Impact of missing entities on missing associations.
The plot supports the assumption that missing an entity promotes missing an association, which the original experiment did not consider and instead attributed the missing associations fully to the use of passive voice femmer2014impact . In fact, the strength of the effect of passive voice on missing associations ($\mu_{RQT}^{PV}=0.38$) is similar to the strength of the effect of missing entities on missing associations ($\mu_{E^{-}}=0.40$). However, the uncertainty of the impact of passive voice ($\sigma_{RQT}^{PV}=0.32$) is higher than that of missing entities ($\sigma_{E^{-}}=0.09$). This means that the effect of missing entities on missing associations is much more reliable than the effect of passive voice on missing associations.
Impact of duration
Figure 11 visualizes the impact of relative duration (i.e., deviation in duration from the overall average time of creating a domain model in minutes) on the two response variables superfluous entities and wrong associations. The red estimate shows that the longer a participant took to generate a domain model (relative duration $>0$), the more likely they were to introduce superfluous entities. The cyan estimate shows that the shorter time a participant took to generate a domain model (relative duration $<0$), the more likely they were to connect an association wrongly.
Figure 11: Impact of relative duration on the number of superfluous entities and wrong associations
Impact of previous training in modeling
Figure 12 visualizes the impact of prior formal training in modeling on the number of wrong associations in a domain model. A participant with prior formal training ($formal=TRUE$) shows a slightly lower likelihood of connecting associations wrongly. The overlapping confidence intervals do, however, indicate a strong variance of the effect.
Figure 12: Marginal effect of prior formal training in modeling on the number of wrong associations
Impact of remaining context factors
None of the remaining context factors has a stronger effect on any of the response variables than the previously mentioned impact visualized in Figure 12. This means that the other context factors are neither notable ($\mu>0.4$) nor significant ($\sigma<\mu$). The replication package contains a detailed summary of all coefficients.
Interaction between domain knowledge and the treatment
Conditional plots visualize interaction effects between two predictors. Figure 13 visualizes the interaction effect between domain knowledge about open source and the treatment on the number of wrong associations. The figure shows that the impact of ambiguous pronouns (cyan whisker plots) on the response variable number of wrong associations diminishes the greater the domain knowledge about open source. For the co-occurrence of ambiguous pronouns and passive voice (purple whisker plots), the effect is less pronounced but symmetrical, i.e., the factor has the strongest impact on the response variable when the domain knowledge is medium. However, the effect contains high uncertainty when the treatment involves ambiguous pronouns, represented by the large and overlapping confidence intervals (cyan and purple whiskers in Figure 13). The collected data does not suffice to support the significance of this effect.
Figure 13: Interaction effect between domain knowledge and the treatment on the number of wrong associations
4.3 Comparison with original study results
Firstly, we compare the results of the original study femmer2014impact with the results of our replication. We compare their missing objects with our merged response variable missing entities, as 16/17 entities in our requirements represent objects, and only one represents an actor. Regarding $H_{0}^{PV\rightarrow E^{-}}$, we obtain conflicting results carver2010towards , as we reject the null hypothesis while the original study does not. We obtain differing results for $H_{0}^{PV\rightarrow A^{-}}$, as we cannot reject it as in the original study.
Secondly, we compare the results of our frequentist data analysis with the results of our Bayesian data analysis. We obtain consistent results carver2010towards for $H_{0}^{RQD\in\{PV,AP,PVAP\}\rightarrow D}$. Neither the frequentist nor Bayesian analysis suggests an impact of the treatment on the relative duration to create a domain model.
[…]
5 Discussion
Section 5.1 answers the research questions. Section 5.2 discusses implications for requirements quality practice and Section 5.3 for requirements quality research. Section 5.4 presents the threats to validity.
5.1 Answers to research questions
5.1.1 Answer to RQ1
RQ1.1: Impact of passive voice.
Using passive voice in natural language requirements specifications has a slight negative effect on the domain modeling activity regarding missing associations. This finding aligns with the conclusions drawn by the original study by Femmer et al. femmer2014impact . However, the Bayesian data analysis emphasizes that both context factors, and especially the number of missing entities, have a significant impact on the number of missing associations as well. Overall, these results support the claim that passive voice can have a negative impact in specific cases but is overall not a significant factor in subsequent activities depending on the requirement femmer2014impact ; krisch2015myth .
RQ1.2: Impact of ambiguous pronouns.
The use of ambiguous pronouns has a strong effect on the number of wrong associations in the resulting domain model. Additionally, using ambiguous pronouns has a slight negative effect on the number of missing and superfluous entities and missing associations. This confirms the risk of using ambiguous pronouns that have been mainly hypothesized in previous research ezzini2022automated and explains the focus on ambiguity in requirements quality research montgomery2022empirical . An ambiguous pronoun in a requirements specification has a 44.6% chance of causing a wrongly connected association in the domain model, limiting the model’s correctness and propagating risk to further activities.
RQ1.3: Combined impact.
The co-occurrence of passive voice and ambiguous pronouns has a strong effect on the number of wrong associations. Additionally, it has a slight effect on the number of missing entities and associations. The impact correlates with but never exceeds the effect of pure, ambiguous pronouns. This supports the assumption that passive voice does not create any further impact in addition to the effect of ambiguous pronouns.
5.1.2 Answer to RQ2
RQ2.1: Impact of context factors.
Only a small number of context factors included in the study show a notable effect on the response variables. The duration of the domain modeling activity confirms assumed patterns: taking shorter than average increases the chance of missing elements or connecting associations wrongly, taking longer time than average increases the chance of adding superfluous entities. Prior formal training in modeling shows a slight yet not significant positive effect.
RQ2.2: Mediation of context factors.
The interaction effect between domain knowledge and the treatment shows that higher domain knowledge can mitigate the negative effect of quality defects on response variables. In particular: higher domain knowledge reduces the chance of connecting associations incorrectly. While the effect still exhibits a large variance, this hints at the possibility of compensating quality defects with domain knowledge.
5.2 Implications for Requirements Quality Practice
The presented results indicate that the negative impact of two requirements quality defects can differ significantly. When allocating resources toward detecting and removing specific quality defects from requirements specifications, organizations can make informed decisions based on the calculated impact of the respective quality factor. In our case, we recommend explicitly detecting and resolving ambiguous pronouns, while passive voice is not critical enough to deserve dedicated attention. This aligns with the common perception in requirements quality research that ambiguity receives the most attention montgomery2022empirical while using passive voice rarely has a tangible impact krisch2015myth . By filtering requirements writing guidelines for quality factors that have a measurable effect, we expect greater acceptance of requirements quality assessment tools in practice franch2020practitioners ; phalp2007assessing .
Additionally, measuring the effect of a quality defect on the relevant attributes of activities that use these requirements allows quantifying it economically frattini2023requirements . While the cost of a quality defect is hard to determine, a company can quantify the cost of activities’ attributes like increased duration. This economic perspective provides additional decision support for companies when assessing whether it is worth detecting and removing a specific quality defect juergens2010much .
Finally, the potential influence of context factors on the impact of quality defects on affected activities may incentivize organizations to invest in developing these factors. For example, improving domain knowledge and providing formal modeling training may compensate quality defects.
5.3 Implications for Requirements Quality Research
Employing Bayesian data analysis to investigate the impact of requirements quality defects provides sophisticated and sensitive insights necessary to propel requirements quality research frattini2023requirements . The result of the analysis models both the direction and strength of an impacting factor while retaining information about its certainty. These insights go beyond the point-wise comparison and binary result of frequentist analyses furia2019bayesian . The frequentist analysis fails to compare the impact of quality defects on the response variables, as even the calculated effect sizes are similar ($0.930.79<|ES|<$ 0.93). The Bayesian data analysis, on the other hand, clearly shows that some effects are much stronger (e.g., $H_{0}^{AP\rightarrow A^{\times}}$) than others (e.g., $H_{0}^{PV\rightarrow A^{-}}$).
Including context factors in the prediction allows comparing the impact of requirements quality with the impact of human and process factors, revealing which causes changes in the response variable. Conducting these studies justifies tool development. Additionally, it increases the likelihood of solving problems with practical relevance franch2020practitioners . However, involving context factors in the evaluation requires accurate metrics. In this study, we followed an approach common in SE research by measuring context factors via self-assessment on a discrete scale. Better metrics of context factors will allow modeling their influence more accurately.
Using a controlled experiment benefits the investigation of the quality factor femmer2014impact . The DAG shown in Figure 9 visualizes this control, as other factors do not influence the treatment in question. This eliminates spurious associations that could confuse the results mcelreath2020statistical . On the other hand, the cost of conducting a controlled experiment—especially with participants from industry—cannot be neglected sjoberg2003challenges . Luckily, statistical causal inference via Bayesian data analysis works equally well with observational data, as shown by Furia et al. furia2023towards .
Worth mentioning is also the synergy between Bayesian data analysis and experiments with a crossover design. The main threats to validity in crossover designs include the impact of confounding variables like the learning and crossover effect kitchenham2003case . The strategies proposed by Vegas et al. vegas2015crossover to mitigate these threats at analysis time are effective yet tedious to implement using frequentist methods. However, within Bayesian data analysis, several of the major threats to validity kitchenham2003case can be addressed by including predictors like the period or the interaction effect between the period and treatment in the model. This eases both the mitigation and evaluation of these threats and naturally factors out the impact of confounding factors caused by the design. Bayesian data analysis promises to be an eligible and accessible tool to evaluate data obtained from crossover-design experiments, harnessing the unique advantages of this design for software engineering research vegas2015crossover .
Finally, Bayesian data analysis allows for incremental improvement of empirical inquiry regarding requirements quality. The causal assumptions that the DAG makes explicit can be reviewed, discussed, and updated to inform future empirical methods. Insights derived from Bayesian data analysis can be used as prior knowledge in subsequent analyses, just as we sensibly used previous results femmer2014impact to inform our priors.
5.4 Threats to Validity
We present and discuss threats that could affect our study based on the guidelines by Wohlin et al. wohlin2012experimentation and extended by the guideline by Vegas et al. vegas2015crossover for the specific threats caused by the use of a crossover design. The threats to validity are prioritized, considering our work focuses on replicating the first study testing a causal theory predicting the impact of requirements quality factors on downstream development tasks.
5.4.1 Internal validity
Our design and the blind nature of the experiment avoid the threat to selection-maturation interaction. Nevertheless, the new settings (i.e., online asynchronous experiment) may have caused a diffusion or imitation of treatments—i.e., information may have been exchanged among the participants. The experiment supervisor monitored the participants to prevent their communication with each other and asked them not to distribute the experimental task and materials. We acknowledge that selection can bias our sample as volunteers are generally more motivated to perform in an experimental task baltes2022sampling .
The crossover design emits additional threats to validity vegas2015crossover . We mitigate the learning by practice effect—i.e., participants getting better when repeating the experimental task—in three ways: Firstly, we disperse the learning effect evenly at design time by randomizing the sequences of treatments. Secondly, we include a warmup object to get participants used to the task and tool but exclude that data from the analysis. Thirdly, we include the period variable as a predictor to factor out the learning effect during the analysis. We avoid the threat of copying by prohibiting communication among participants and using experimental objects where solutions cannot be copied from one task to another.
The threat of optimal sequence describes the risk that there is a sequence in which the treatment is applied, which optimizes the participants’ performance in deriving domain models. We cannot block this threat at analysis time as the sequence and participant IDs are highly correlated. This is because we could—in all but one cases—assign only one participant ($n_{p}=25$) to each sequence ($n_{s}=n_{r}!=4!=24$). Because of this strong correlation, the Bayesian model is incapable of distinguishing between the impact of the sequence ($\beta_{seq}$) from the within-participant variance ($\alpha_{PID}$vegas2015crossover . More participants per sequence would have been necessary to block the threat of an optimal sequence, but these were unavailable to us.
Finally, we address the threat of carryover—i.e., the change of the impact caused by the period in which the treatment was applied—at analysis time by including the term $period*treatment$ in the predictors. This way, the carryover effect is factored out from the impact of the treatment and analyzable from the posterior distributions.
5.4.2 Conclusion validity
We addressed the reliability of measures threat by peer-reviewing the extraction of the dependent variables from the collected domain models, for which we report an inter-rater agreement of 87.23%. We addressed the random heterogeneity of the subjects by a design in which each participant acts as their own control group. Moreover, we focused on including and analyzing context variables related to the participants’ experience. Our sample of participants is not representative of all context factors. Consequently, our Bayesian data analysis cannot identify all causal effects of some context factors. However, by including them in the causal considerations, the effect of the factors is isolated from the potential confounding variables leven2022broken . The conclusion validity of our study is strengthened by applying two different data analysis approaches and comparing their results. The data analysis suffers from the threat of low statistical power when it comes to evaluating interaction effects, as reliably identifying them requires a larger sample size gelman2018interaction . We limit the number of interaction effects considered in our models and discuss the uncertainty around the coefficient estimates to minimize this threat.
5.4.3 Construct validity
Our study can suffer from mono-operation bias as we focus only on a subset of quality factors that can potentially exist montgomery2022empirical ; frattini2022live . Nevertheless, our goal with this replication is to extend the initial quality factor of passive voice reported by Femmer et al. femmer2014impact to a second one—ambiguous pronouns—which is widespread as indicated by the literature frattini2022live ; montgomery2022empirical . Similarly, a confounding of constructs and level of constructs could influence the outcomes of our study. For example, the presence of several ambiguous pronouns or passive voice sentences rather than their binary presence or absence from a specification. Further replications, focusing on improving construct validity, should include several levels of each treatment. Mono-method bias is a potential threat to construct validity—i.e., we measured the dependent variables using a single type of measurement, following the original study. However, the measurements were based on a pre-defined protocol and peer-reviewed. Our study may result in a restricted generalizability across constructs since the presence or absence of the different quality factors could result in side effects for other interesting outcomes we did not measure (e.g., comprehensibility or maintainability of the specification). Among the social threats to construct validity, we acknowledge that hypothesis guessing may have taken place since the participants could try to guess the concrete goal of the experimental task based on the invitation text and material provided during the sessions. Nevertheless, we used the same text and phrasing to invite all participants and the same material during the experimental task. Finally, evaluation apprehension could have played a role since some participants are students at the authors’ institution. However, students did not receive rewards (e.g., extra grade points) for participating in the experiment, and they received their course grades before the start of the experiment.
5.4.4 External validity
The main threat to the external validity of this study is the interaction of setting and treatment as the size and the complexity of the selected specifications, despite being sampled from a real-world data set, might not be representative of the industrial practice.
6 Conclusion
Requirements quality research lacks empirical evidence and research strategies to advance beyond proposing and following normative rules with unclear impact frattini2022live to better understanding and solving problems relevant to practice franch2017practitioners ; franch2020practitioners . In the scope of our study, we replicated the only controlled experiment on the impact of requirements quality defects on subsequent activities. We demonstrated a method of evaluating data collected through a controlled experiment using a crossover design with Bayesian data analysis. We showed the impact (1) of requirements quality defects varies and (2) may be mediated by context and confounding factors.
We can confidently support the recommendation of SE researchers to adopt Bayesian data analysis to improve causal reasoning and inference furia2019bayesian ; furia2022applying ; torkar2020missing , which will propel requirements quality research. This shift requires focusing on problems such as scrutinizing the explicit causal assumptions of a DAG, visualizing requirements quality impact, evolving prior knowledge about their impact, and comparing models concerning their predictive power.
We envision that adopting sophisticated statistical tools like Bayesian data analysis and the focus of empirical studies on investigating the impact of requirements quality defects will steer requirements quality research in a relevant and effective direction. We hope that the documentation of this study inspires fellow researchers to adopt our method and tools for replication.
Acknowledgements.
This work was supported by the KKS foundation through the S.E.R.T. Research Profile project at Blekinge Institute of Technology. We are deeply grateful to Parisa Yousefi from Ericsson AB for recruiting practitioners to the experiment.
Conflict of interest
The authors declare that they have no conflict of interest.
Data Availability Statement
All supplementary material, including protocols and guidelines for data collection and extraction, the raw data, analysis scripts, figures, and results, are available in our replication package. The replication package is available on GitHub at https://github.com/JulianFrattini/rqi-proto and archived on Zenodo at https://doi.org/10.5281/zenodo.10423666.